can someone help me I have a question similar to this and I am completely lost. How do you determine the between vs within groups?

Three sets of three mice were randomly selected to be placed in a standard maze but with different color doors. The response is the time required to complete the maze as seen in the table below. Perform appropriate analysis to test whether there is an effect due to door coloring.

Color

Time

Yi.

ȳi.

ȳ..

Red

9

11

10

30

10

11.89

Green

20

21

17

58

19.33

Black

6

5

8

19

6.33

H0:                                                                                                                   HA:

ANOVA Table

Source

Df

SS

MS

F

p

Between groups

Within groups

Total
Answer #1

The summary statistics obtained from the given data are as below.

No Music Soft Rock Classical
Total 30 58 19
n 3 3 3
Mean 10.00 19.33 6.33
Sum Of Squares 2 8.67 4.67

The Hypothesis:

H0: There is no difference between the means of the timings for the three colored doors.

Ha: The mean of at least one time is different from the others..

________________________________________________

The ANOVA table is as below. the p value is calculated for F = 52.97 for df1 = 2 and df2 = 6

The Fcritical is calculated at = 0.05 (default level) for df1 = 2 and df2 = 6

Source DF SS Mean Square F Fcv p
Between 2 269.55 134.78 52.97 5.143 0.0002
Within 6 15.27 2.54
Total 8 284.82

The Decision Rule:

If Ftest is > F critical, Then Reject H0.

Also if p-value is < , Then reject H0.

The Decision:

Since Ftest (52.97) is > F critical (5.143), We Reject H0.

Also since p-value (0.0002) is < (0.05, default), We Reject H0.

The Conclusion: There is sufficient evidence at the 95% level of significance to conclude that the mean of at least one time is different from the others.

__________________________________________________________

Calculations For the ANOVA Table:

Overall Mean = Average of Means = (10 + 19.33 + 6.33) / 3 = 11.89

SS treatment = SUM [n* ( - overall mean)2] = 3 * (10 - 11.89)2 + 3 * (19.33 - 11.89)2 + 3 * (6.33 - 11.89)2 = 269.55

df1 = k - 1 = 3 - 1 = 2

MSTR = SS treatment/df1 = 269.55 / 2 = 134.78

SSWithin = SUM (Sum of Squares) = 2 + 8.67 + 4.67 = 15.27

df2 = N - k = 9 - 3 = 6

Therefore MS within = SSwithin/df2 = 15.27 / 6 = 2.54

F = MSTR/MSE = 134.78 / 2.54 = 52.97

______________________________________________

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image